### v2

# Imperfect Competition in Markets for Short-Circuit Current Services

Abstract
An important limitation of Inverter-Based Resources (IBR) is their reduced contribution to Short-Circuit Current (SCC), as compared to that of Synchronous Generators (SGs). With increasing penetration of IBR in most power systems, the reducing SCC poses challenges to a secure system operation, as line protections may not trip when required. In order to address this issue, the SCC ancillary service could be procured via an economic mechanism, aiming at securing adequate SCC on all buses. However, the suitability of markets for SCC services is not well understood, given that these could be prone to market-power issues: since the SCC contributions from various SGs to a certain bus are determined by the electrical topology of the grid, this is a highly local service. It is necessary to understand if SGs at advantageous electrical locations could exert market power and, if so, how it could be mitigated. In order to fill this gap, this paper adopts an SCC-constrained bilevel model to investigate strategic behaviors of SGs. To address the non-convexity due to unit commitment variables, the model is restructured through a primal-dual formulation. Based on a modified IEEE 30-bus system, cases with strategic SGs placed at different buses are analyzed. These studies demonstrate that agents exerting market power could achieve up to triple revenues from SCC provision, highlighting the need to carefully design these markets.
keywords:
Ancillary Services , Short-Circuit Current , Market Power , Bilevel Optimization , Primal-Dual Formulation
journal: Sustainable Energy, Grids and Networks
Nomenclature
Abbreviations and Acronyms
  • IBR
    Inverter-Based Resources
  • KKT
    Karush–Kuhn–Tucker
  • $L$
    Lower-Level
  • SCC
    Short-Circuit Current
  • SG
    Synchronous Generator
  • UC
    Unit Commitment
  • UL
    Upper-Level
Indices and Sets
  • $b,\mathcal{B}$
    Index, Set of buses
  • $c,\mathcal{C}$
    Index, Set of IBR
  • $g,\mathcal{G}$
    Index, Set of generic SGs
  • $\hat{g},\hat{\mathcal{G}}$
    Index, Set of strategic SGs
  • $\check{g},\check{\mathcal{G}}$
    Index, Set of non-strategic SGs
  • $m,\mathcal{M}$
    Index, Set of commitment decision pairs of SGs
  • $t,T$
    Index, Set of time periods
  • $\Psi(g),\Phi(c)$
    Index of bus for SGs and IBR
Constants and Parameters
  • $\upalpha_{c,t}$
    Capacity factor of IBR at time period $t$
  • $\overline{\upbeta}^{\textrm{m}}_{\hat{g}}$
    Upper bound of the strategic bidding multiplier for SGs in energy market
  • $\overline{\upbeta}^{\textrm{SCC}}_{\hat{g},b}$
    Upper bound of the strategic bidding multiplier for SGs in SCC service market of bus $b$
  • $\textrm{I}_{b_{\textrm{lim}}}$
    Minimum requirement of SCC for bus $b$ (p.u.)
  • $\textrm{K}^{\textrm{st}}_{g},\textrm{K}^{\textrm{sh}}_{g}$
    Start-up and shut-down costs of SGs (€/h)
  • $\textrm{k}_{bg},\textrm{k}_{bc},\textrm{k}_{bm}$
    Approximate coefficients of the SCC on bus $b$ (p.u.)
  • $\textrm{O}_{g}^{\textrm{nl}}$
    No-load costs of SGs (€/h)
  • $\textrm{O}_{g,b,t}^{\textrm{SCC}}$
    Commitment price of SGs for providing SCC to bus $b$ at time period $t$ (€/h)
  • $\textrm{O}_{g}^{\textrm{m}}$
    Marginal power generation costs of SGs (€/MWh)
  • $\textrm{O}_{c}^{\textrm{l}}$
    Energy bids of IBR (€/MWh)
  • $\textrm{P}^{\textrm{D}}_{t}$
    Total system demand at time period $t$ (MW)
  • $\textrm{P}_{g}^{\textrm{min}},\textrm{P}_{g}^{\textrm{max}}$
    Minimum stable generation and rated power of SGs (MW)
  • $\textrm{P}_{g}^{\textrm{rd}},\textrm{P}_{g}^{\textrm{ru}}$
    Ramp-down and ramp-up limits of SGs (MW/h)
  • $\textrm{P}^{\textrm{max}}_{c}$
    Rated power of IBR (MW)
  • W
    Penalty factor in the primal-dual formulation
Primal variables
  • $\beta_{\hat{g},t}^{\textrm{m}}$
    Bidding multiplier of strategic SGs in energy market at time period $t$
  • $\beta_{\hat{g},b,t}^{\textrm{SCC}}$
    Bidding multiplier of strategic SGs in SCC service market of bus $b$ at time period $t$
  • $C_{g,t}^{\textrm{st}},C_{g,t}^{\textrm{sh}}$
    Start-up/shut-down costs incurred by SGs at time period $t$ (€/h)
  • $P_{g,t}$
    Power output of SGs at time period $t$ (MW)
  • $P_{c,t}$
    Power output of IBR at time period $t$ (MW)
  • $u_{g,t}$
    Binary variable, commitment of SG at time period $t$
Dual variables
  • $\lambda^{\textrm{E}}_{t}$
    Energy clearing price at time period $t$ (€/MWh)
  • $\lambda_{b,t}^{\textrm{SCC}}$
    SCC clearing price for bus $b$ at time period $t$ (€/p.u.)
  • ${\mu}^{\textrm{min}}_{g,t},{\mu}^{\textrm{max}}_{g,t}$
    Associated with the generation limits of SGs
  • $\pi_{g,t}^{\textrm{rd}},\pi_{g,t}^{\textrm{ru}}$
    Associated with the ramp constraints of SGs
  • $\rho_{g,t}^{\textrm{st}},\sigma_{g,t}^{\textrm{st}},\rho_{g,t}^{\textrm{sh}},\sigma_{g,t}^{\textrm{sh}}$
    Associated with the constraints for start-up/shut-down costs of SGs
  • $\zeta_{c,t}^{\textrm{min}},\zeta_{c,t}^{\textrm{max}}$
    Associated with the constraints for generation limits of IBR
  • $\psi_{g,t}^{\textrm{min}},\psi_{g,t}^{\textrm{max}}$
    Associated with the constraints for relaxation of commitment status of SGs
1 Introduction
Eelectric power systems throughout the world are shifting to renewable energy dominated architectures for achieving the goal of net-zero carbon emissions in coming decades. This requires the massive installation of Inverter-Based Resources (IBR), mainly wind and solar generators.
Since power electronic inverters have very limited overload capacity, they only contribute to about one fifth of the Short-Circuit Current (SCC) provided by Synchronous Generators (SGs) [1, 2]. Sufficient SCC is a main criterion for a secure system operation [3, 4], as lack of this physical quantity can have important negative impacts, including causing protection devices to fail in responding to a short-circuit fault. Deficiency in SCC may also lead to voltage oscillations, potentially causing the phase-locked loop to lose synchronization.
Various proposals have been made to overcome these issues in IBR-dominated systems, including inserting phase-change material into chips to improve the overload capability of grid inverters [5], and installing synchronous condensers to maintain the minimum SCC of critical buses [6]. Another feasible solution is considering SCC as an ancillary service, through dispatching SGs to offer sufficient SCC during certain periods, thus guaranteeing an adequate performance of protections during short-circuit faults [7]. For power systems with both SGs and IBR, a methodology for evaluating SCC on specific buses within dispatch models was developed in [8]. It computes the SCC contributions according to the Unit Commitment (UC) status and equivalent impedance between thermal units and buses concerned. It is important to note that online SGs provide little SCC to distant buses, and only contribute significantly to increasing SCC in their surrounding nodes. Therefore, once the SCC requirement of all buses is satisfied through dispatching certain critically-located SGs, the service provided by other ones will no longer be needed. This highly local nature of the SCC service may give certain generation companies some room to exert market power for increasing their own profits, particularly when they manage SGs located at advantageous electrical locations.
Market power issues in the electricity industry have been extensively analyzed. Authors in [9] investigated if market power is one of the reasons for the price increases in the wholesale electricity market of India. Other regions have also been examined, including California [10, 11], England and Wales [12], and the Nordic wholesale markets [13]. The relationship between market power, congestion, and renewable energy has also been explored, such as in the Italian day-ahead market [14]. Market power mitigation mechanisms in electricity markets have been proposed [15], including applying price caps as an efficient way to mitigate market power in terms of transmission congestion [16]. These efforts, among many others, demonstrate the importance of understanding if certain agents may exert market power in electricity systems.
Regarding modeling techniques for capturing imperfect competition effects, bilevel optimization has proven to be a powerful tool. Through a bilevel model one can simulate a situation where a leader makes decisions before the followers, who then react to those decisions [17]. Leaders, or price makers, are the agents with the capability to exert market power, due for example to their larger size in comparison to others, or their privileged location within a network. Mathematically, bilevel models are generally solved through replacing lower-level problems (representing the followers) by their equivalent Karush-Kuhn-Tucker (KKT) conditions, which are then included as constraints in upper level, therefore leading to an equivalent single-level model [17]. However, this approach requires the lower-level problem to be continuous and convex, as otherwise the KKT conditions either do not exactly correspond to the original problem, or simply can not be derived.
Several techniques for tackling non-convex bilevel models have been proposed. An iterative methodology based on column-and-constraint generation for the original problem [18] has been widely employed. In particular, non-convex bilevel problems for energy markets are investigated in [19, 20, 21], where [19, 20] address the challenge of commitment variables in UC, while [21] tackles the binary variables that flag the installation of phase shifting transformers. Binary variables appear in both levels in [22] to characterize the interaction and on/off states of different objectives. Reference [23] also includes binary variables, in this case to consider the installation of certain equipment. However, this approach first used in [18] would lead to a bilevel model independent of bids of strategic players, which contradicts our purpose of capturing potential strategic behaviors. An alternative method is to derive KKT conditions for each enumerated values of all binary variables [24]. This has the disadvantage of increasing the number of KKT conditions exponentially, especially when it comes to a large-scale problem with more combinations of binary states being considered, leading to high computational complexity. Therefore, the approach used in the present paper for solving the SCC-constrained bilevel UC problem is the one introduced in [25]. It is based on a primal-dual formulation which can tackle the non-convexity in the lower level, and being suitable for our purposes as the strategic bidding decisions for SCC may be considered as optimized variables.
In this work, we start by describing SCC constraints and verifying their effectiveness in a linear form. Then, for each network at hand, buses that face potential short-circuit risk are identified, and marginal prices for SCC services offered by various SGs are computed. Finally, the UC bilevel problem representing the market is solved via a primal-dual formulation, allowing to analyze the impact of strategic operation of different thermal units on the market outcome.
Specifically, the novel contributions of this work are:
  1. 1.
    For the first time, a market formulation considering SCC as a service is proposed, by means of a bilevel optimization model. The non-convexity in the problem, due to the binary commitment state of SGs, is addressed via a primal-dual reformulation.
  2. 2.
    The proposed model allows to provide the first analysis of imperfect competition in markets for SCC services. We demonstrate that strategic SGs could achieve up to triple revenues from SCC markets, highlighting the need to carefully design these markets.
The remainder of this paper is structured as follows: Section 2 describes modeling of the SCC constraint and its application in the bilevel model. Section 3 explains how to tackle the non-convexity of the model by primal-dual reformulation. Section 4 includes relevant case studies to analyze market power issues, while Section 5 gives the conclusion and suggests lines for future work.
2 Short-Circuit Current Constrained Bilevel Problem Modeling
This section introduces the SCC constraints that are next applied into a bilevel model to secure all buses to an acceptable level of SCC through a market setting. Note that only three-phase nodal short-circuit faults are considered. For the sake of clarity, time index ‘$t$’ is omitted in the SCC expressions.
2.1 Formulation of Short-Circuit Current Constraints
The SCC constraints derived in [8] are used in this work, which incorporate the current contributions from both SGs and IBR. For a power system with bus $b\in\mathcal{B}$, SGs $g\in\mathcal{G}$ and IBR $c\in\mathcal{C}$, the SCC on bus $b$ can be expressed as:
$I_{b_{\textrm{SC}}}=\frac{\sum_{g\in\mathcal{G}}Z_{b\Psi(g)}\textrm{I}_{g}u_{g}+\sum_{c\in\mathcal{C}}Z_{b\Phi(c)}\textrm{I}_{c}\upalpha_{c}}{Z_{bb}}$ (1)
where (1) accounts for the discrete nature of SCC from SGs through their commitment variable $u_{g}$. Furthermore, the SCC contribution is also dependent on the impedance $Z_{ij}$ and, for the case of IBR, their capacity factor $\upalpha_{c}$. $\textrm{I}_{g}$ and $\textrm{I}_{c}$ are the short-circuit injections from SGs and IBR, respectively.
The impedance matrix is calculated by inverting the corresponding admittance matrix, $Z_{ij}$ is thus hard to include in a duality-involved optimization problem. Therefore, a data-driven method is adopted to approximate the actual SCC constraints [8]. Coefficients $\{k_{b_{g}},k_{b_{c}},k_{b_{m}}\}$ are introduced to approximate the actual values of SCC (1) into:
$I_{b_{\textrm{L}}}=\sum_{g}k_{bg}u_{g}+\sum_{c}k_{bc}\upalpha_{c}+\sum_{m}k_{bm}\eta_{m}\geq\textrm{I}_{b_{\textrm{lim}}}$ (2a)
$\eta_{m}=u_{g_{\textrm{1}}}u_{g_{\textrm{2}}},\quad\textrm{s.t.}~\{g_{\textrm{1}},g_{\textrm{2}}\}=m$ (2b)
$m\in\mathcal{M}=\{g_{\textrm{1}},g_{\textrm{2}}\mid\forall g_{\textrm{1}},\forall g_{\textrm{2}}\in\mathcal{G}\}$ (2c)
where (2a) is the approximate value of SCC, which is forced to be higher than $\textrm{I}_{b_{\textrm{lim}}}$. $\eta_{m}$ captures the interactions between pairs of SGs to simulate the nonlinearity, defined as (2b)-(2c). The approximate coefficients $k_{(\cdot)}$ are determined by operating an optimization-based classification procedure, comprising of a data generation process and a minimization problem that tries to reduce the error led by the approximation. This classification only serves to train the coefficients in (2) for fitting the actual SCC, meaning that it is in fact an offline process to the bilevel optimization model. An example for the training process used in this work can be accessed in [26].
2.2 Bilevel Model with Short-Circuit Current Constraints
The coefficients $\{k_{b_{g}},k_{b_{c}},k_{b_{m}}\}$ in (2a) are constants in the bilevel model as discussed above, therefore italics are no longer used for them. Besides, the number of dual variables corresponding to the equations for linearizing the quadratic term $\eta_{m}$ increases exponentially with more SGs involved, leading to high computational complexity. Therefore, $\eta_{m}$ is removed from the SCC constraint in the following models, still yielding an effective formulation as demonstrated in Section 4.2.
2.2.1 Upper-Level Model
Strategic SGs maximize the profits at the Upper-Level (UL) problem, where the optimal bidding decisions are determined to form the bids reported to the lower level. The UL model is expressed as:
$\displaystyle\max_{V_{\textrm{UL}}}\sum_{t}\Biggr[\sum_{\hat{g}}\biggl(\lambda^{\textrm{E}}_{t}P_{\hat{g},t}+\sum_{b}\lambda_{b,t}^{\textrm{SCC}}\textrm{k}_{b\hat{g}}u_{\hat{g},t}$
$\qquad-\textrm{O}_{\hat{g}}^{\textrm{nl}}u_{\hat{g},t}-\textrm{O}_{\hat{g}}^{\textrm{m}}P_{\hat{g},t}-C^{\textrm{st}}_{\hat{g},t}-C^{\textrm{sh}}_{\hat{g},t}\biggl)\Biggr]$ (3a)
where:
$V_{\textrm{UL}}=\{\beta_{\hat{g},t}^{\textrm{m}},\beta_{\hat{g},b,t}^{\textrm{SCC}}\}$ (3b)
subject to:
$1\leq\beta_{\hat{g},t}^{\textrm{m}}\leq\overline{\upbeta}^{\textrm{m}}_{\hat{g}}\quad\forall\hat{g},\forall t$ (3c)
$1\leq\beta_{\hat{g},b,t}^{\textrm{SCC}}\leq\overline{\upbeta}^{\textrm{SCC}}_{\hat{g},b}\quad\forall\hat{g},\forall b,\forall t$ (3d)
where objective function (3a) indicates that the revenues of strategic SGs come from the provision of energy and SCC, meanwhile the costs include no-load costs, marginal generation costs, start-up and shut-down costs. Formula (3b) shows the variables in UL problem. Constraints (3c)-(3d) confine the strategic multipliers of bids.
2.2.2 Lower-Level Model
The system operator in the Lower-Level (LL) problem takes the strategic bids decided by (3b), then minimizes the system-wide operation costs. The LL model is defined as:
$\displaystyle\min_{V_{\textrm{LL}}}\sum_{t}\Biggr[\sum_{\hat{g}}\biggl(\textrm{O}_{\hat{g}}^{\textrm{nl}}u_{\hat{g},t}+\sum_{b}\textrm{O}_{\hat{g},b,t}^{\textrm{SCC}}\beta_{\hat{g},b,t}^{\textrm{SCC}}u_{\hat{g},t}+\textrm{O}_{\hat{g}}^{\textrm{m}}\beta_{\hat{g},t}^{\textrm{m}}P_{\hat{g},t}$
$\qquad+C_{\hat{g},t}^{\textrm{st}}+C_{\hat{g},t}^{\textrm{sh}}\biggl)+\sum_{\check{g}}\biggl(\textrm{O}_{\check{g}}^{\textrm{nl}}u_{\check{g},t}+\sum_{b}\textrm{O}_{\check{g},b,t}^{\textrm{SCC}}u_{\check{g},t}$
$\qquad+\textrm{O}_{\check{g}}^{\textrm{m}}P_{\check{g},t}+C_{\check{g},t}^{\textrm{st}}+C_{\check{g},t}^{\textrm{sh}}\biggl)+\sum_{c}\textrm{O}_{c}^{\textrm{l}}P_{c,t}\Biggr]$ (4a)
where:
$V_{\textrm{LL}}=\Bigl\{u_{\hat{g},t},P_{\hat{g},t},C_{\hat{g},t}^{\textrm{st}},C_{\hat{g},t}^{\textrm{sh}},$
$\hskip 5.69046pt\qquad\quad u_{\check{g},t},P_{\check{g},t},C_{\check{g},t}^{\textrm{st}},C_{\check{g},t}^{\textrm{sh}},P_{c,t}\Bigl\}$ (4b)
subject to:
$\sum_{\hat{g}}\textrm{k}_{b\hat{g}}u_{\hat{g},t}+\sum_{\check{g}}\textrm{k}_{b\check{g}}u_{\check{g},t}$
$+\sum_{c}\textrm{k}_{bc}\upalpha_{c,t}\geq\textrm{I}_{b_{\textrm{lim}}}:(\lambda_{b,t}^{\textrm{SCC}})\quad\forall b,\forall\hat{g},\forall\check{g},\forall c,\forall t$ (4c)
$\textrm{P}^{\textrm{D}}_{t}=\hskip-3.0pt\sum_{\hat{g}}\hskip-3.0ptP_{\hat{g},t}+\sum_{\check{g}}\hskip-3.0ptP_{\check{g},t}+\sum_{c}\hskip-3.0ptP_{c,t}:(\lambda_{t}^{\textrm{E}})\quad\forall\hat{g},\forall\check{g},\forall c,\forall t$ (4d)
$u_{\hat{g},t}\textrm{P}_{\hat{g}}^{\textrm{min}}\leq P_{\hat{g},t}\leq u_{\hat{g},t}\textrm{P}_{\hat{g}}^{\textrm{max}}:({\mu}^{\textrm{min}}_{\hat{g},t},{\mu}^{\textrm{max}}_{\hat{g},t})\quad\forall\hat{g},\forall t$ (4e)
$u_{\check{g},t}\textrm{P}_{\check{g}}^{\textrm{min}}\leq P_{\check{g},t}\leq u_{\check{g},t}\textrm{P}_{\check{g}}^{\textrm{max}}:({\mu}^{\textrm{min}}_{\check{g},t},{\mu}^{\textrm{max}}_{\check{g},t})\quad\forall\check{g},\forall t$ (4f)
$-\textrm{P}_{\hat{g}}^{\textrm{rd}}\leq P_{\hat{g},t}-P_{\hat{g},(t-1)}\leq\textrm{P}_{\hat{g}}^{\textrm{ru}}:({\pi}^{\textrm{rd}}_{\hat{g},t},{\pi}^{\textrm{ru}}_{\hat{g},t})\quad\forall\hat{g},\forall t$ (4g)
$-\textrm{P}_{\check{g}}^{\textrm{rd}}\leq P_{\check{g},t}-P_{\check{g},(t-1)}\leq\textrm{P}_{\check{g}}^{\textrm{ru}}:({\pi}^{\textrm{rd}}_{\check{g},t},{\pi}^{\textrm{ru}}_{\check{g},t})\quad\forall\check{g},\forall t$ (4h)
$C_{\hat{g},t}^{\textrm{st}}\geq 0:(\rho_{\hat{g},t}^{\textrm{st}})\quad\forall\hat{g},\forall t$ (4i)
$C_{\hat{g},t}^{\textrm{sh}}\geq 0:(\rho_{\hat{g},t}^{\textrm{sh}})\quad\forall\hat{g},\forall t$ (4j)
$C_{\check{g},t}^{\textrm{st}}\geq 0:(\rho_{\check{g},t}^{\textrm{st}})\quad\forall\check{g},\forall t$ (4k)
$C_{\check{g},t}^{\textrm{sh}}\geq 0:(\rho_{\check{g},t}^{\textrm{sh}})\quad\forall\check{g},\forall t$ (4l)
$C_{\hat{g},t}^{\textrm{st}}\geq(u_{\hat{g},t}-u_{\hat{g},(t-1)})\textrm{K}^{\textrm{st}}_{\hat{g}}:(\sigma_{\hat{g},t}^{\textrm{st}})\quad\forall\hat{g},\forall t$ (4m)
$C_{\hat{g},t}^{\textrm{sh}}\geq(u_{\hat{g},(t-1)}-u_{\hat{g},t})\textrm{K}^{\textrm{sh}}_{\hat{g}}:(\sigma_{\hat{g},t}^{\textrm{sh}})\quad\forall\hat{g},\forall t$ (4n)
$C_{\check{g},t}^{\textrm{st}}\geq(u_{\check{g},t}-u_{\check{g},(t-1)})\textrm{K}^{\textrm{st}}_{\check{g}}:(\sigma_{\check{g},t}^{\textrm{st}})\quad\forall\check{g},\forall t$ (4o)
$C_{\check{g},t}^{\textrm{sh}}\geq(u_{\check{g},(t-1)}-u_{\check{g},t})\textrm{K}^{\textrm{sh}}_{\check{g}}:(\sigma_{\check{g},t}^{\textrm{sh}})\quad\forall\check{g},\forall t$ (4p)
$0\leq P_{c,t}\leq\upalpha_{c,t}\textrm{P}^{\textrm{max}}_{c}:(\zeta_{c,t}^{\textrm{min}},\zeta_{c,t}^{\textrm{max}})\quad\forall c,\forall t$ (4q)
$u_{\hat{g},t},u_{\check{g},t}\in\{0,1\}\quad\forall\hat{g},\forall\check{g},\forall t$ (4r)
where the system operating costs are represented in (4a), including the costs explained in (3a), the costs of SCC service and energy provided by IBR. (4b) includes primal variables in the LL model which follow constraints: SCC on bus $b$ contributed by SGs and IBR (4c); Demand-supply power balance constraints (4d); Minimum stable generation and maximum generation constraints for SGs (4e)-(4f); Ramp-up/down constraints (4g)-(4h); Start-up/shut-down costs (4i)-(4p); Minimum and maximum generation constraints for IBR (4q); (4r) impose integrality for commitment variables of SGs.
We compute the SCC service offer $\textrm{O}_{g,b,t}^{\textrm{SCC}}$ via a marginal unit pricing [7], from which the economic value of SCC from SG $g$ for bus $b$, i.e., $\textrm{k}_{bg}$, is obtained as the commitment price (this indicates that the service offer is implicitly expressed as $\textrm{O}_{g,b,t}^{\textrm{SCC}}/\textrm{k}_{bg}$, with the unit being €/p.u.). This computation implies obtaining the difference in objective values between two problems. First, problem (4) is solved as an energy-only UC without SCC-related terms in the objective. Second, the SCC provision $\textrm{k}_{bg}$ is eliminated in (4c), then calculating hourly operation costs of the modified dispatch problem. The difference in those two solutions represents the cost of forcing other units to startup for providing SCC. This whole process is further explained through a numerical example in Section 4.3.
Furthermore, since the commitment of SGs has an impact in energy provision and SCC contribution, these two services are highly coupled. As a result, some SCC contributions come as by-products of power balance or securing certain buses rather than given ones, thereby being essentially free (i.e., $\textrm{O}_{g,b,t}^{\textrm{SCC}}$ is zero). In summary, the value of this two-step procedure for computing $\textrm{O}_{g,b,t}^{\textrm{SCC}}$ is to recognize which SCC contributions to critical buses come with non-zero service costs.
3 Solution Methodology
Given that the bilevel model in this work contains binary variables in the lower level through (4r), converting the LL model into the equivalent KKT conditions becomes impractical. A primal-dual formulation is adopted to tackle this non-convex bilevel model, which is designed to restate the LL problem by deriving the dual with relaxed binary variables. Then, after recovering the discreteness of commitment decisions, the bilevel problem can be further converted to a single-level modeling by minimizing the duality gap between the primal and dual, as we discuss next.
3.1 Dualization of Primal Lower-Level Problem
The binary variables are relaxed to be continuous within a limited range, indicated as follows:
$0\leq u_{\hat{g},t}\leq 1:(\psi_{\hat{g},t}^{\textrm{min}},\psi_{\hat{g},t}^{\textrm{max}})\quad\forall\hat{g},\forall t$ (5a)
$0\leq u_{\check{g},t}\leq 1:(\psi_{\check{g},t}^{\textrm{min}},\psi_{\check{g},t}^{\textrm{max}})\quad\forall\check{g},\forall t$ (5b)
with this relaxation, the dual variables for all constraints in (4) can then be defined. Consequently, the primal LL problem is now defined by (4a)-(4q) and (5).
After the relaxation, the dual problem associated with the relaxed LL problem is derived as follows:
$\displaystyle\max_{V_{\textrm{DLL}}}\sum_{t}\Biggr[\textrm{P}^{\textrm{D}}_{t}\lambda_{t}^{\textrm{E}}+\sum_{b}\textrm{I}_{b_{\textrm{lim}}}\lambda_{b,t}^{\text{SCC}}-\hskip-7.11317pt\sum_{c}\upalpha_{c,t}\textrm{P}^{\textrm{max}}_{c}\zeta_{c,t}^{\textrm{max}}-\hskip-7.11317pt\sum_{\hat{g}}\psi_{\hat{g},t}^{\textrm{max}}$
$-\sum_{\check{g}}\psi_{\check{g},t}^{\textrm{max}}\Biggr]+\sum_{\hat{g},(t=1)}\hskip-7.11317ptP_{\hat{g},0}\pi_{\hat{g},t}^{\textrm{rd}}-\sum_{\hat{g},t}\textrm{P}_{\hat{g}}^{\textrm{rd}}\pi_{\hat{g},t}^{\textrm{rd}}+\sum_{\check{g},(t=1)}\hskip-7.11317ptP_{\check{g},0}\pi_{\check{g},t}^{\textrm{rd}}$
$-\sum_{\check{g},t}\textrm{P}_{\check{g}}^{\textrm{rd}}\pi_{\check{g},t}^{\textrm{rd}}-\sum_{\hat{g},(t=1)}\hskip-7.11317ptP_{\hat{g},0}\pi_{\hat{g},t}^{\textrm{ru}}-\sum_{\hat{g},t}\textrm{P}_{\hat{g}}^{\textrm{ru}}\pi_{\hat{g},t}^{\textrm{ru}}-\sum_{\check{g},(t=1)}\hskip-7.11317ptP_{\check{g},0}\pi_{\check{g},t}^{\textrm{ru}}$
$-\sum_{\check{g},t}\textrm{P}_{\check{g}}^{\textrm{ru}}\pi_{\check{g},t}^{\textrm{ru}}-\sum_{\hat{g},(t=1)}\hskip-7.11317ptu_{\hat{g},0}\textrm{K}_{\hat{g}}^{\textrm{st}}\sigma_{\hat{g},t}^{\textrm{st}}-\sum_{\check{g},(t=1)}\hskip-7.11317ptu_{\check{g},0}\textrm{K}_{\check{g}}^{\textrm{st}}\sigma_{\check{g},t}^{\textrm{st}}$
$+\sum_{\hat{g},(t=1)}\hskip-7.11317ptu_{\hat{g},0}\textrm{K}_{\hat{g}}^{\textrm{sh}}\sigma_{\hat{g},t}^{\textrm{sh}}+\sum_{\check{g},(t=1)}\hskip-7.11317ptu_{\check{g},0}\textrm{K}_{\check{g}}^{\textrm{sh}}\sigma_{\check{g},t}^{\textrm{sh}}$ (6a)
where:
$V_{\textrm{DLL}}=\Bigl\{\lambda_{t}^{\textrm{E}},\lambda_{b,t}^{\textrm{SCC}},\zeta_{c,t}^{\textrm{max}},$
$\qquad\qquad\mu^{\textrm{min}}_{\hat{g},t},{\mu}^{\textrm{max}}_{\hat{g},t},{\pi}^{\textrm{rd}}_{\hat{g},t},{\pi}^{\textrm{ru}}_{\hat{g},t},\sigma_{\hat{g},t}^{\textrm{st}},\sigma_{\hat{g},t}^{\textrm{sh}},\psi_{\hat{g},t}^{\textrm{max}},$
$\qquad\qquad\mu^{\textrm{min}}_{\check{g},t},{\mu}^{\textrm{max}}_{\check{g},t},{\pi}^{\textrm{rd}}_{\check{g},t},{\pi}^{\textrm{ru}}_{\check{g},t},\sigma_{\check{g},t}^{\textrm{st}},\sigma_{\check{g},t}^{\textrm{sh}},\psi_{\check{g},t}^{\textrm{max}}\Bigl\}$ (6b)
subject to:
$\textrm{O}_{\hat{g}}^{\textrm{nl}}+\sum_{b}\beta_{\hat{g},b,t}^{\textrm{SCC}}\textrm{O}_{\hat{g},b,t}^{\textrm{SCC}}-\textrm{k}_{b\hat{g}}\lambda_{b,t}^{\textrm{SCC}}-\textrm{P}_{\hat{g}}^{\textrm{max}}\mu_{\hat{g},t}^{\textrm{max}}+\textrm{P}_{\hat{g}}^{\textrm{min}}\mu_{\hat{g},t}^{\textrm{min}}$
$+\textrm{K}_{\hat{g}}^{\textrm{st}}(\sigma_{\hat{g},t}^{\textrm{st}}-\sigma_{\hat{g},(t+1)}^{\textrm{st}})+\textrm{K}_{\hat{g}}^{\textrm{sh}}(\sigma_{\hat{g},(t+1)}^{\textrm{sh}}-\sigma_{\hat{g},t}^{\textrm{sh}})$
$+\psi_{\hat{g},t}^{\textrm{max}}\geq 0,\quad\forall\hat{g},\forall b,\forall t\leq T-1$ (6c)
$\textrm{O}_{\hat{g}}^{\textrm{nl}}+\sum_{b}\beta_{\hat{g},b,t}^{\textrm{SCC}}\textrm{O}_{\hat{g},b,t}^{\textrm{SCC}}-\textrm{k}_{b\hat{g}}\lambda_{b,t}^{\textrm{SCC}}-\textrm{P}_{\hat{g}}^{\textrm{max}}\mu_{\hat{g},t}^{\textrm{max}}+\textrm{P}_{\hat{g}}^{\textrm{min}}\mu_{\hat{g},t}^{\textrm{min}}$
$+\textrm{K}_{\hat{g}}^{\textrm{st}}\sigma_{\hat{g},t}^{\textrm{st}}-\textrm{K}_{\hat{g}}^{\textrm{sh}}\sigma_{\hat{g},t}^{\textrm{sh}}+\psi_{\hat{g},t}^{\textrm{max}}\geq 0,\quad\forall\hat{g},\forall b,\forall t=T$ (6d)
$\textrm{O}_{\check{g}}^{\textrm{nl}}+\sum_{b}\textrm{O}_{\check{g},b,t}^{\textrm{SCC}}-\textrm{k}_{b\check{g}}\lambda_{b,t}^{\textrm{SCC}}-\textrm{P}_{\check{g}}^{\textrm{max}}\mu_{\check{g},t}^{\textrm{max}}+\textrm{P}_{\check{g}}^{\textrm{min}}\mu_{\check{g},t}^{\textrm{min}}$
$+\textrm{K}_{\check{g}}^{\textrm{st}}(\sigma_{\check{g},t}^{\textrm{st}}-\sigma_{\check{g},(t+1)}^{\textrm{st}})+\textrm{K}_{\check{g}}^{\textrm{sh}}(\sigma_{\check{g},(t+1)}^{\textrm{sh}}-\sigma_{\check{g},t}^{\textrm{sh}})$
$+\psi_{\check{g},t}^{\textrm{max}}\geq 0,\quad\forall\check{g},\forall b,\forall t\leq T-1$ (6e)
$\textrm{O}_{\check{g}}^{\textrm{nl}}+\sum_{b}\textrm{O}_{\check{g},b,t}^{\textrm{SCC}}-\textrm{k}_{b\check{g}}\lambda_{b,t}^{\textrm{SCC}}-\textrm{P}_{\check{g}}^{\textrm{max}}\mu_{\check{g},t}^{\textrm{max}}+\textrm{P}_{\check{g}}^{\textrm{min}}\mu_{\check{g},t}^{\textrm{min}}$
$+\textrm{K}_{\check{g}}^{\textrm{st}}\sigma_{\check{g},t}^{\textrm{st}}-\textrm{K}_{\check{g}}^{\textrm{sh}}\sigma_{\check{g},t}^{\textrm{sh}}+\psi_{\check{g},t}^{\textrm{max}}\geq 0,\quad\forall\check{g},\forall b,\forall t=T$ (6f)
$\textrm{O}_{\hat{g}}^{\textrm{m}}b_{\hat{g},t}^{\textrm{m}}-\lambda_{t}^{\textrm{E}}+\mu^{\textrm{max}}_{\hat{g},t}-\mu^{\textrm{min}}_{\hat{g},t}+\pi^{\textrm{ru}}_{\hat{g},t}-\pi^{\textrm{ru}}_{\hat{g},(t+1)}$
$-\pi^{\textrm{rd}}_{\hat{g},t}+\pi^{\textrm{rd}}_{\hat{g},(t+1)}\geq 0,\quad\forall\hat{g},\forall t\leq T-1$ (6g)
$\textrm{O}_{\hat{g}}^{\textrm{m}}b_{\hat{g},t}^{\textrm{m}}-\lambda_{t}^{\textrm{E}}+\mu^{\textrm{max}}_{\hat{g},t}-\mu^{\textrm{min}}_{\hat{g},t}+\pi^{\textrm{ru}}_{\hat{g},t}-\pi^{\textrm{rd}}_{\hat{g},t}\geq 0,$
$\quad\forall\hat{g},\forall t=T$ (6h)
$\textrm{O}_{\check{g}}^{\textrm{m}}-\lambda_{t}^{\textrm{E}}+\mu^{\textrm{max}}_{\check{g},t}-\mu^{\textrm{min}}_{\check{g},t}+\pi^{\textrm{ru}}_{\check{g},t}-\pi^{\textrm{ru}}_{\check{g},(t+1)}$
$-\pi^{\textrm{rd}}_{\check{g},t}+\pi^{\textrm{rd}}_{\check{g},(t+1)}\geq 0,\quad\forall{\check{g}},\forall t\leq T-1$ (6i)
$\textrm{O}_{\check{g}}^{\textrm{m}}-\lambda_{t}^{\textrm{E}}+\mu^{\textrm{max}}_{\check{g},t}-\mu^{\textrm{min}}_{\check{g},t}+\pi^{\textrm{ru}}_{\check{g},t}-\pi^{\textrm{rd}}_{\check{g},t}\geq 0,$
$\quad\forall\check{g},\forall t=T$ (6j)
$1-\sigma_{\hat{g},t}^{\textrm{st}}\geq 0,\quad\forall\hat{g},\forall t$ (6k)
$1-\sigma_{\hat{g},t}^{\textrm{sh}}\geq 0,\quad\forall\hat{g},\forall t$ (6l)
$1-\sigma_{\check{g},t}^{\textrm{st}}\geq 0,\quad\forall\check{g},\forall t$ (6m)
$1-\sigma_{\check{g},t}^{\textrm{sh}}\geq 0,\quad\forall\check{g},\forall t$ (6n)
$\textrm{O}_{c}^{\textrm{l}}-\lambda_{t}^{\textrm{E}}+\zeta^{\textrm{max}}_{c,t}\geq 0,\quad\forall c,\forall t$ (6o)
$V_{\textrm{DLL}}\in\mathbb{R}_{+},\quad\forall b,\forall\hat{g},\forall\check{g},\forall c,\forall t$ (6p)
where (6a) is the objective function of the dual problem. (6b) shows the dual variables involved. Constraints (6c)-(6f), (6g)-(6j), (6k)-(6n) and (6o) represent dual constraints regarding primal variables: $u_{\hat{g},t},u_{\check{g},t}$; $P_{\hat{g},t},P_{\check{g},t}$; $C_{\hat{g},t}^{\textrm{st}},C_{\hat{g},t}^{\textrm{sh}},C_{\check{g},t}^{\textrm{st}},C_{\check{g},t}^{\textrm{sh}}$ and $P_{c,t}$, respectively. Constraints (6p) express non-negativity of the dual variables. Note that $\lambda_{t}^{\textrm{E}}$ is also lower bounded by zero here to ensure non-negative energy revenues for SGs.
3.2 Primal-Dual Single-Level Optimization Model
Binary variables are only relaxed for deriving the dual problem. By recovering (4r), they are constrained to be integer again to capture the UC property. This implies that the reformulated LL model is not completely equivalent to the primal, thus holding weak duality. The penalty function method with a parameter W can be employed to minimize the Duality Gap (DG) to produce feasible solutions, as introduced in [25]. Therefore, the final form of the problem is:
[…]
where:
[…]
$V=\{V_{\textrm{UL}},V_{\textrm{LL}},V_{\textrm{DLL}}\}$ (7c)
subject to:
[…]
where (7) is a single-level problem. Function (7a) couples the maximization of UL objective (3a) and minimization of DG between primal and dual LL problems. The DG is defined as (7b). Formulas (7c) and (7d) clarify the variables and constraints involved. The positive penalty factor W governs the trade-offs of objectives. An appropriate value for this parameter must be found for the problem at hand, as we discuss in Section 4.4, for which we define a ratio associated with primal and dual LL problems, to evaluate the error of solutions as follows:
[…]
where a small value of $r^{\textrm{DG}}$ indicates a minor gap between the primal and dual problems, i.e., a more accurate solution.
It is noteworthy that problem (7) is still nonlinear, as it still contains quadratic terms $\lambda^{\textrm{E}}_{t}P_{\hat{g},t}$, $\lambda_{b,t}^{\textrm{SCC}}\textrm{k}_{b\hat{g}}u_{\hat{g},t}$ in (3a), $\textrm{O}_{\hat{g},b,t}^{\textrm{SCC}}\beta_{\hat{g},b,t}^{\textrm{SCC}}u_{\hat{g},t}$, $\textrm{O}_{\hat{g}}^{\textrm{m}}\beta_{\hat{g},t}^{\textrm{m}}P_{\hat{g},t}$ in (4a). McCormick envelopes are employed to perform the linearization with proper upper bounds on relevant variables, as the setting in [25]. The envelopes for those second-order terms are shown as follows.
Envelope 1: The term $\lambda^{\textrm{E}}_{t}P_{\hat{g},t}$ that follows $0\leq\lambda^{\textrm{E}}_{t}\leq\overline{\lambda}^{\textrm{E}}$ can be reformulated as:
$\lambda_{t}^{\textrm{E}}\textrm{P}_{\hat{g}}^{\textrm{min}}\leq z^{{\textrm{Re}^{\textrm{E}}}}_{\hat{g},t}\leq\overline{\lambda}^{\textrm{E}}P_{\hat{g},t}+\textrm{P}_{\hat{g}}^{\textrm{min}}(\lambda_{t}^{\textrm{E}}-\overline{\lambda}^{\textrm{E}})$ (9a)
$\overline{\lambda}^{\textrm{E}}P_{\hat{g},t}+\textrm{P}_{\hat{g}}^{\textrm{max}}(\lambda_{t}^{\textrm{E}}-\overline{\lambda}^{\textrm{E}})\leq z^{{\textrm{Re}^{\textrm{E}}}}_{\hat{g},t}\leq\lambda_{t}^{\textrm{E}}\textrm{P}_{\hat{g}}^{\textrm{max}}$ (9b)
where $z^{{\textrm{Re}^{\textrm{E}}}}_{\hat{g},t}=\lambda^{\textrm{E}}_{t}P_{\hat{g},t}$. The upper bound should consider marginal costs that may be reported by both strategic and non-strategic units. Therefore, $\overline{\lambda}^{\textrm{E}}=\mathrm{max}~\{\textrm{O}_{\hat{g}}^{\textrm{m}}\overline{\upbeta}^{\textrm{m}}_{\hat{g}},\textrm{O}_{\check{g}}^{\textrm{m}}\mid\forall\hat{g},\forall\check{g}\}$.
Envelope 2: For term $\lambda_{b,t}^{\textrm{SCC}}\textrm{k}_{b\hat{g}}u_{\hat{g},t}$ complying with constraint $0\leq\lambda_{b,t}^{\textrm{SCC}}\leq\overline{\lambda}_{b,t}^{\textrm{SCC}}$, it is restated as:
$0\leq z^{{\textrm{Re}^{\textrm{SCC}}}}_{\hat{g},b,t}\leq\overline{\lambda}_{b,t}^{\textrm{SCC}}\textrm{k}_{b\hat{g}}u_{\hat{g},t}$ (10a)
$\lambda_{b,t}^{\textrm{SCC}}\textrm{k}_{b\hat{g}}-\overline{\lambda}_{b,t}^{\textrm{SCC}}\textrm{k}_{b\hat{g}}(1-u_{\hat{g},t})\leq z^{{\textrm{Re}^{\textrm{SCC}}}}_{\hat{g},b,t}\leq\lambda_{b,t}^{\textrm{SCC}}\textrm{k}_{b\hat{g}}$ (10b)
where $z^{{\textrm{Re}^{\textrm{SCC}}}}_{\hat{g},b,t}=\lambda_{b,t}^{\textrm{SCC}}\textrm{k}_{b\hat{g}}u_{\hat{g},t}$. Similarly, the upper limit is set as $\overline{\lambda}_{b,t}^{\textrm{SCC}}=\mathrm{max}~\{\overline{\upbeta}^{\textrm{SCC}}_{\hat{g},b}\textrm{O}_{\hat{g},b,t}^{\textrm{SCC}}/\textrm{k}_{b\hat{g}},\textrm{O}_{\check{g},b,t}^{\textrm{SCC}}/\textrm{k}_{b\check{g}}\mid\forall\hat{g},\forall\check{g}\}$.
Envelope 3: The nonlinear term $\textrm{O}_{\hat{g},b,t}^{\textrm{SCC}}\beta_{\hat{g},b,t}^{\textrm{SCC}}u_{\hat{g},t}$, constrained by (3d), is remodeled as:
$z^{{\textrm{Bid}^{\textrm{SCC}}}}_{\hat{g},b,t}\geq\textrm{O}_{\hat{g},b,t}^{\textrm{SCC}}u_{\hat{g},t}$ (11a)
$z^{{\textrm{Bid}^{\textrm{SCC}}}}_{\hat{g},b,t}\geq\textrm{O}_{\hat{g},b,t}^{\textrm{SCC}}(\overline{\upbeta}^{\textrm{SCC}}_{\hat{g},b}(u_{\hat{g},t}-1)+\beta_{\hat{g},b,t}^{\textrm{SCC}})$ (11b)
$z^{{\textrm{Bid}^{\textrm{SCC}}}}_{\hat{g},b,t}\leq\textrm{O}_{\hat{g},b,t}^{\textrm{SCC}}\overline{\upbeta}^{\textrm{SCC}}_{\hat{g},b}u_{\hat{g},t}$ (11c)
$z^{{\textrm{Bid}^{\textrm{SCC}}}}_{\hat{g},b,t}\leq\textrm{O}_{\hat{g},b,t}^{\textrm{SCC}}(u_{\hat{g},t}+\beta_{\hat{g},b,t}^{\textrm{SCC}}-1)$ (11d)
where $z^{{\textrm{Bid}^{\textrm{SCC}}}}_{\hat{g},b,t}=\textrm{O}_{\hat{g},b,t}^{\textrm{SCC}}\beta_{\hat{g},b,t}^{\textrm{SCC}}u_{\hat{g},t}$.
Envelope 4: Likewise, term $\textrm{O}_{\hat{g}}^{\textrm{m}}\beta_{\hat{g},t}^{\textrm{m}}P_{\hat{g},t}$, following (3c), is derived as:
$z^{{\textrm{Bid}^{\textrm{E}}}}_{\hat{g},t}\geq\textrm{O}_{\hat{g}}^{\textrm{m}}(P_{\hat{g},t}+\textrm{P}_{\hat{g}}^{\textrm{min}}(\beta_{\hat{g},t}^{\textrm{m}}-1))$ (12a)
$z^{{\textrm{Bid}^{\textrm{E}}}}_{\hat{g},t}\geq\textrm{O}_{\hat{g}}^{\textrm{m}}(\overline{\upbeta}^{\textrm{m}}_{\hat{g}}P_{\hat{g},t}+\textrm{P}_{\hat{g}}^{\textrm{max}}(\beta_{\hat{g},t}^{\textrm{m}}-\overline{\upbeta}^{\textrm{m}}_{\hat{g}}))$ (12b)
$z^{{\textrm{Bid}^{\textrm{E}}}}_{\hat{g},t}\leq\textrm{O}_{\hat{g}}^{\textrm{m}}(\overline{\upbeta}^{\textrm{m}}_{\hat{g}}P_{\hat{g},t}+\textrm{P}_{\hat{g}}^{\textrm{min}}(\beta_{\hat{g},t}^{\textrm{m}}-\overline{\upbeta}^{\textrm{m}}_{\hat{g}}))$ (12c)
$z^{{\textrm{Bid}^{\textrm{E}}}}_{\hat{g},t}\leq\textrm{O}_{\hat{g}}^{\textrm{m}}(P_{\hat{g},t}+\textrm{P}_{\hat{g}}^{\textrm{max}}(\beta_{\hat{g},t}^{\textrm{m}}-1))$ (12d)
where $z^{{\textrm{Bid}^{\textrm{E}}}}_{\hat{g},t}=\textrm{O}_{\hat{g}}^{\textrm{m}}\beta_{\hat{g},t}^{\textrm{m}}P_{\hat{g},t}$.
In summary, to solve the nonlinear bilevel model with non-convexity, binary variables are relaxed to be continuous to derive the dual of primal LL problem. Then they are tightened to achieve realistic UC solutions, although this leads to a DG between the primal and dual LL problems. The penalty function method is used to minimize the DG, giving a trade-off between the solution accuracy and values of objectives. Finally, nonlinear terms are restated to form the final single-level, mixed-integer linear program.
4 Case Studies
4.1 Test System Setting
Figure 1: Modified IEEE 30-bus system.
Table 1: Operating Parameters of Synchronous Generators
Bus 2 3 4 5 27 30
$\textrm{O}_{g}^{\textrm{nl}}$ (€/h) 1,743 1,501 1,376 1,093 990 857
$\textrm{O}_{g_{1}}^{\textrm{m}}$ (€/MWh) 6.20 7.10 10.47 12.28 13.53 15.36
$\textrm{O}_{g_{2}}^{\textrm{m}}$ (€/MWh) 7.07 8.72 11.49 12.84 14.60 15.02
$\textrm{K}_{g}^{\textrm{st}}$ (€/h) 20,000 12,500 9,250 7,200 5,500 3,100
$\textrm{K}_{g}^{\textrm{sh}}$ (€/h) 5,000 2,850 1,850 1,440 1,200 1,000
$\textrm{P}_{g}^{\textrm{min}}$ (MW) 658 576 302 133 130 58
$\textrm{P}_{g}^{\textrm{max}}$ (MW) 1,317 1,152 756 667 650 576
$\textrm{P}_{g}^{\textrm{rd}}$ (MW/h) 263 230 302 267 390 346
$\textrm{P}_{g}^{\textrm{ru}}$ (MW/h) 263 230 302 267 390 346
$u_{g,0}$ 1 1 1 1 1 0
$P_{g,0}$ (MW) 1,054 922 605 534 520 0
Case studies on a modified IEEE 30-bus system are carried out in this work. The diagram of the test system is shown in Fig. 1. The IBR, i.e., wind turbines are added to buses {1, 23, 26}, with energy bids {2.46, 3.19, 2.86} €/MWh, respectively. In order to model a modular SCC contribution from SGs, we consider that two units are located in each bus containing an SG, i.e., buses {2, 3, 4, 5, 27, 30}. This way, diverse decisions can be made in the market. For example, different UC status (i.e., both online, both offline, just one of them online) can determine various SCC contributions and bids. The SGs in bus 2 are separately named as $g_{1}$-bb2 and $g_{2}$-bb2, with 2gg-bb2 standing for both of them. Other thermal units follow the same nomenclature. The SGs’ parameters are given in Table 1.
The network impedances and reactances of SGs refer to [28] and [29], respectively. Load demand $\textrm{P}^{\textrm{D}}_{t}\in[\textrm{5109},\textrm{7689}]$ MW. Total installed wind generation capacity is 750 MW. Base power is $S_{B}=\textrm{100}~\textrm{MVA}$. $\textrm{I}_{c}$, the short-circuit injection from IBR, is 1.00 p.u. Nominal voltage is set as 0.95 p.u. for conservativeness [8]. Maximum multipliers for bidding decisions $\overline{\upbeta}^{\textrm{m}}_{\hat{g}}$ and $\overline{\upbeta}^{\textrm{SCC}}_{\hat{g},b}$ are both set as 2. All code and data are available at the public repository [26]. The simulations in this work were processed by Julia-JuMP and Gurobi in version 12.0.0.
4.2 Validation of Short-Circuit Current Constraints
The approximate SCC constraints (2a) are linearized to be (4c) for the primal-dual formulation. Two evaluations are employed to validate those SCC constraints in the first place.
Evaluation 1: Training errors of approximation. Two types of errors are analyzed to verify the approximation [7]: Type-I: $I_{b_{\textrm{L}}}\geq\textrm{I}_{b_{\textrm{lim}}}\cap I_{{b_{\textrm{SC}}}}<\textrm{I}_{b_{\textrm{lim}}}$ means the approximate SCC is looser than the actual requirement, a potentially dangerous situation; Type-II: $I_{b_{\textrm{L}}}<\textrm{I}_{b_{\textrm{lim}}}\cap I_{{b_{\textrm{SC}}}}\geq\textrm{I}_{b_{\textrm{lim}}}$ implies that approximate SCC defines a tighter boundary than the grid-code value $\textrm{I}_{b_{\textrm{lim}}}$, therefore satisfying system security. Type-I errors are likely to cause insecure operating conditions, thus being undesired.
Evaluation 2: Minimum SCC distribution. We consider all buses to comply with approximate and linear SCC constraints respectively in an economic dispatch problem. The minimum SCC of all buses can then be calculated to verify if each of them at any time periods is beyond $\textrm{I}_{b_{\textrm{lim}}}$.
Table 2: Assessment of Minimum SCC and Training Errors for Different SCC Requirements
$\textrm{I}_{b_{\textrm{lim}}}$ (p.u.) Minimum SCC Type-I Type-II
Approximate (2a) Linearized (4c) $N_{e}$ errerr $N_{e}$ errerr
p.u. Bus p.u. Bus
3.00 4.50 30 4.73 30 0 0 2 -1.13%
4.00 4.13 30 4.06 30 0 0 8 -0.56%
5.00 5.04 26 5.20 30 0 0 11 -1.79%
Results evaluated at different $\textrm{I}_{b_{\textrm{lim}}}$ are presented in Table 2, in which $N_{e}$ and errerr respectively indicate total number and averaged value of approximation errors. The approximation eliminates Type-I errors and produces only a small amount of Type-II errors, indicating a good performance. The minimum SCC of the entire system is above the $\textrm{I}_{b_{\textrm{lim}}}$, meaning that the SCC level on all buses are always within the safe range, justifying the approximate SCC constraints (2a). The effectiveness of the linearized SCC constraint (4c) is demonstrated by the results on the middle section of Table 2, showing that minimum SCC is still beyond $\textrm{I}_{b_{\textrm{lim}}}$.
4.3 System Operation under Perfect Competition
Here, we examine a perfectly competitive case with a complete operation cycle (i.e., 24 hours as the length of market horizon), in which all SGs are assumed to submit true offers, so the bilevel model reduces to a single-level problem as (4). This will serve as a model to compute honest SCC bids for various SGs and identify critical buses that need to follow SCC constraints.
First, we remove SCC-related terms from the model to simulate an energy-only market, i.e., with no guarantees on adequate SCC levels. From the solution of this problem, the resulting minimum SCC for each bus is obtained. Any SCC lower than $\textrm{I}_{b_{\textrm{lim}}}$ means that the corresponding bus is at risk and should in fact incorporate an explicit SCC constraint. After adding these SCC constraints, using a value of $\textrm{I}_{b_{\textrm{lim}}}=\textrm{5}~\textrm{p.u.}$ in all system buses as in [8], truthful bids for this service are calculated through the marginal unit pricing introduced in [7], as explained in Section 2.2.
Figure 2: Minimum SCC levels in all buses with and without explicit constraints.
Figure 3: Marginal service prices for units providing SCC to critical buses in one-day operation. Prices in other cases are always zero, therefore not shown.
The SCC distribution without explicit constraints is illustrated in Fig. 2, confirming minimum SCC in buses {26, 29, 30} to be below the acceptable threshold. In this market setting, only units on buses {2, 3} are always committed to supply energy due to their low marginal costs, providing SCC meanwhile. It is worth noting that the SCC of bus 1 is beyond $\textrm{I}_{b_{\textrm{lim}}}$, though its local IBR can only provide up to 1 p.u. of SCC. That is because the bus is electrically close to online SGs, therefore benefiting from a safe level of SCC directly as a by-product of the energy market. However, the SCC on buses {26, 29, 30} may be insufficient within the day as they are at the edge of the system, being away from the typically committed thermal units (surrounding SGs are not always dispatched due to their high marginal costs).
The marginal SCC prices payable to different units for critical buses {26, 29, 30} are depicted in Fig. 3, in which the prices for bus 30 are in fact the same for several units (blue line in right plot) due to their very similar SCC contribution to this bus. For the case of bus 26 (left plot), only SGs in bus 27 would capture prices for SCC provision, whereas other SGs are not chosen to provide this service. This is due to the fact that 2gg-bb27 are closer to bus 26, providing a level of SCC which is already sufficient to secure that bus, while other units provide a lower volume, resulting in those SCC contributions being less cost-effective. This explains that the elimination of those excess contributions do not incur higher costs. Vice versa, without SCC provision from $g_{1}$-bb27 or $g_{2}$-bb27, another local unit or more distant units with limited SCC contribution would have to be committed for meeting the security requirement, rendering higher costs which lead to these marginal SCC prices for units in bus 27.
For bus 30 (right plot), more than half of SGs participate in this service market capturing a certain price, which is generally lower than that for the case of bus 26 since most units provide low SCC due to long electrical distances, implying that the service from any of them can be replaced by any other. The marginal service price for bus 29 is zero, meaning that it shows sufficient SCC once other buses are secured, so no extra units would be required.
In summary, from this perfectly competitive market case it can be observed that only units $g_{1}$-bb27 and $g_{2}$-bb27 provide SCC to bus 26 at non-zero prices, while the market of bus 30 has multiple units at different electrical locations providing the service. Consequently, given the test system and operation data adopted, thermal units that do not offer SCC services to critical buses are naturally not profitable in the SCC market. The revenue term $\lambda_{b,t}^{\textrm{SCC}}\textrm{k}_{bg}u_{g,t}$’ is thus eliminated from their objective function in following studies.
4.4 System Operation under Imperfect Competition
The bilevel model introduced in Section 2.2 allows to represent the case where a power company may behave strategically in the market to increase profits. We assume this company to operate several SGs, so that market power issues may be observed through analyzing different cases:
Table 3: Evaluation of the Impact of Penalty Factor on Strategic Cases (Units of $r^{\textrm{DG}}$’ and ‘Profits’ are % and k€/day, respectively)
Bus of strategic units
W Bus 2 Bus 3 Bus 4 Bus 5 Bus 27
$r^{\textrm{DG}}$ Pro $r^{\textrm{DG}}$ Pro $r^{\textrm{DG}}$ Pro $r^{\textrm{DG}}$ Pro $r^{\textrm{DG}}$ Pro
1 3.60 423 4.07 341 6.79 155 7.24 146 7.39 292
[…]
100 -0.38 262 0.87 208 3.59 13 4.10 -3 2.78 103
1000 -0.38 261 0.87 197 3.59 12 4.10 -4 2.78 103
Table 4: Averaged Strategic Multipliers of SGs in the Energy Market
Strategic units Bus 2 Bus 3 Bus 4 Bus 5 Bus 27
$\hat{g}_{1}$ $\hat{g}_{2}$ $\hat{g}_{1}$ $\hat{g}_{2}$ $\hat{g}_{1}$ $\hat{g}_{2}$ $\hat{g}_{1}$ $\hat{g}_{2}$ $\hat{g}_{1}$ $\hat{g}_{2}$
$\sum_{t}\beta_{\hat{g},t}^{\textrm{m}}/T$ 1.71 1.67 1.54 1.29 1.05 1.03 1.00 1.00 1.04 1.00
Figure 4: Shadow prices of SCC constraints for bus 26 and bus 30 when two strategic SGs are placed in a single bus. The price for bus 29 is zero, so not graphically shown.
Table 5: Revenues of SGs for Providing SCC and Corresponding Service Payment (k€/day)
[…]
Bus 2 Bus 3 Bus 4 Bus 5 Bus 27
[…]
4.4.1 Two Strategic SGs in a Single Bus
The strategic company in this setting owns two SGs placed in a single bus, thus producing five scenarios, i.e., buses {2, 3, 4, 5, 27} (bus 30 is not considered, as the units placed there are too expensive to be dispatched to provide services). The $r^{\textrm{DG}}$ (8) and profits (accounting for both energy and SCC markets) of the strategic company in all five cases are given in Table 3, with values for the penalty term W of {1, 10, 100, 1000}. As the penalty factor increases, $r^{\textrm{DG}}$ and the profits decrease. It is noteworthy that $r^{\textrm{DG}}$ could be negative, as the primal LL problem is relaxed to be linear, thus having a bound potentially lower than the dual [27]. From these results, we set going forward that the strategic firm gives equal weight to the DG (i.e., $\textrm{W}=\textrm{10}$) in all scenarios, since a noticeably small $r^{\textrm{DG}}$ can be achieved while yielding a substantial profit, the ultimate objective of the strategic company. Although higher profits could be obtained with the setting $\textrm{W}=\textrm{1}$, it actually suggests that the strategic firm can not make a reasonable evaluation regarding its profitability [25].
While market power from SCC services is the focus of this work, we still briefly analyze market power in the energy clearing, by checking averaged bidding decisions, i.e., $\sum_{t}\beta_{\hat{g},t}^{\textrm{m}}/T$, shown in Table 4. Units at buses {2, 3} are seen to strategically bid to maximize profits, whereas other expensive units normally offer true energy bids even if they are capable of misreporting them (note that a bidding multiplier of 1 indicates a truthful bid). This is because there are significantly cheaper units with important capacity in the energy market, making high-cost units difficult to be dispatched. Consequently, the optimal solution for them is bidding honestly in order to clear the market. Market power in SCC services is analyzed next.
[…]
These results demonstrate that market power may be enhanced through occupying advantageous electrical positions, as these allow to provide SCC exclusively for critical buses, therefore achieving a greater manipulation of service prices.
Payment and revenues for SCC services: The revenues of SGs in various strategic scenarios are specified in Table 5. This table shows revenues for generic SGs (first column) under different strategic SG pairs (right columns), with bold diagonal elements representing strategic SGs’ revenues, and other values indicating non-strategic SGs’ income.
[…]
In summary, to increase revenues from SCC markets, certain strategic SGs could both manipulate service prices and commit for longer periods, further exacerbating the market distortion. This has been shown to be particularly relevant for the critical location in terms of SCC in this test network, bus 26, where the SCC market is monopolized by the electrically-closest generators placed in bus 27.
4.4.2 Two Strategic SGs in Different Buses
Figure 6: Comparison of SCC clearing prices for bus 26 and bus 30 when two strategic SGs are placed at different buses.
We now analyze the impact of a strategic company owning two generators placed in different system buses, instead of co-locating them in the same bus. Given that there exist many combinations of strategic SGs placed at different locations, we properly select only two thermal units to carry out case studies. In this subsection, the scenario of strategic $\hat{g}_{1}$-bb2 & $\hat{g}_{2}$-bb27’ is considered, as it represents general principles that other scenarios also follow. The penalty factor W is still set to 10, with $r^{\textrm{DG}}=\textrm{1.02\%}$.
[…]
These results show that it is best for the strategic company to concentrate their generators around the most critical location in terms of SCC requirements, which in this test network is bus 26. If the company were to make an investment in a second SG, assuming they already own one placed in bus 27 (the electrically closest generator bus to the critical bus 26), it shows to be a better decision to build it in the same bus 27 than splitting their generation assets geographically: indeed, case $\hat{g}_{1}$-bb2 & $\hat{g}_{2}$-bb27’ reduces the company’s capability to manipulate SCC prices.
[…]
In summary, market power issues exacerbate when the number of strategic units in the same critical location increases, rather than a company owning generators in different buses. This suggests that procuring SCC by other means, e.g., by investing in a synchronous compensator owned by the Transmission System Operator (TSO) in that critical location, could significantly mitigate market power. Then, the SCC market may be maintained for the rest of the system, as it shows to be fairly competitive in the test network considered.
5 Conclusion and Future Work
A bilevel model has been introduced to explore potential market power issues in SCC services, allowing to capture strategic behaviors of some agents. Given the non-convex nature of the lower-level problem, as it contains binary commitment variables for SGs, directly replacing it by its KKT conditions is not a suitable approach for eventually obtaining an equivalent single-level problem. Therefore, a primal-dual formulation is employed instead to solve the model, which has been shown to work effectively in this setting.
Case studies based on a modified IEEE 30-bus system have revealed how strategic units exert market power in SCC services to achieve higher revenues. Generators located next to critical buses in terms of SCC requirements have been shown to effectively manipulate prices for this service in their own benefit. Furthermore, they could also stay online for a longer time, in order to both capture the higher prices for a sustained period and create barriers for other providers of this service. This has the double effect of limiting the profits of non-strategic units and leading to a higher service cost to consumers, highlighting the need to carefully design SCC markets.
Given the issues identified with regards to imperfect competition in SCC markets, the main line for future work is to develop strategies to mitigate this market power. Potential countermeasures could be considered at both the market and technical levels, such as limiting the bid or clearing price of SCC services, weighing the negative impacts of installing new units in buses that are prone to trigger market power when planning the system, and investing in the nonprofitable synchronous condenser to a proper location.